# Semilinear spaces of matrices with transitivity, rank or spectral constraints

### v1

@@ 6.3 The full structure theorem for optimal spaces
Let $V$ be a (right) $\mathbb{D}$-vector space with finite dimension $n>0$. Assume that $|\mathbb{F}|\geq nd$, and let $b$ be a nonisotropic $\mathbb{F}$-bilinear form on $V$ such that $\mathcal{S}:=\mathcal{A}_{b,\mathbb{D}}$ is an optimal irreducible $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$.
We shall prove that $\dim\operatorname{\operatorname{Alt}}(\mathcal{S})=1$ in any case.
  • Assume first that $n\geq\max(4-d,2)$. Then $\dim\mathcal{S}=\alpha(n,d)\geq\alpha(n,d)-n+\max(4-d,2)$ and the claimed statement is deduced from point (c) of Theorem 1.20.
  • Assume now that $n=2$ and $d=1$. Then $\dim\mathcal{S}=1$, and we take $u_{0}\in\mathcal{S}\smallsetminus\{0\}$. Since $\mathcal{S}$ is irreducible, $u_{0}$ must be invertible, and we deduce that $\dim\mathcal{S}x=1=nd-1$ for all $x\in V\smallsetminus\{0\}$. Moreover $\mathcal{S}$ is target-reduced because $u_{0}$ is surjective. Hence Lemma 5.3 shows that $\dim\operatorname{\operatorname{Alt}}(\mathcal{S})\leq 1$, and we conclude that $\dim\operatorname{\operatorname{Alt}}(\mathcal{S})=1$ because $b\in\operatorname{\operatorname{Alt}}(\mathcal{S})$.
  • If $n=1$ and $d>1$ then $\dim_{\mathbb{F}}(\mathcal{S}x)=\dim\mathcal{S}=nd-1$ for all $x\in V\smallsetminus\{0\}$, and obviously $\mathcal{S}$ is target-reduced, so once more we find by Proposition 5.3 that $\dim\operatorname{\operatorname{Alt}}(\mathcal{S})=1$.
  • If $n=1$ and $d=1$ then $\mathcal{S}=\{0\}$ and $\operatorname{\operatorname{Alt}}(\mathcal{S})$ is the space of all $\mathbb{F}$-bilinear forms on $V$, which has dimension $1$.
Then, we apply Proposition 5.2 and find that $\mathbb{D}$ has quadratic type over $\mathbb{F}$ and that, for an associated pair $(\sigma,e)$, there exists a $\sigma$-sesquilinear form $B:V^{2}\rightarrow\mathbb{D}$ such that $b:(x,y)\mapsto e(B(x,y))$. Besides, it has already been observed in Section 2.1 that $B$ is uniquely determined by $b$.
@@ 6.4 The quadratic type case
We finish this section by proving the results that are specific to the division algebras of quadratic type: Theorems 1.10 and 1.11.
So, assume that $\mathbb{D}$ has quadratic type over $\mathbb{F}$, and denote by $(\sigma,e)$ an associated pair. Let $V$ be an $n$-dimensional vector space over $\mathbb{D}$, with $n\geq 1$. Let $\mathcal{S}$ be an irreducible optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ such that $|\mathbb{F}|\geq nd$.
If $n\geq 2$ then we apply Theorems 1.7 and 1.8 to find a $\sigma$-sesquilinear nondegenerate form $B:V^{2}\rightarrow\mathbb{D}$ such that $\mathcal{S}=\mathcal{A}_{b,\mathbb{D}}$ for $b:(x,y)\mapsto e(B(x,y))$.
Assume now that $n=1$. In that case we need only consider the special case where $V=\mathbb{D}$. By the isomorphism $a\in\mathbb{D}\mapsto[x\mapsto ax]\in\operatorname{\operatorname{End}}_{\mathbb{D}}(\mathbb{D})$ of $\mathbb{F}$-vector spaces, we associate to $\mathcal{S}$ an $\mathbb{F}$-linear hyperplane $H$ of $\mathbb{D}$. Thus $H$ is the kernel of $x\mapsto e(\alpha x)$ for some $\alpha\in\mathbb{D}^{\times}$. Then, for all $a\in H$, we have
@@
This shows that the nondegenerate bilinear form $b:(x,y)\in\mathbb{D}^{2}\mapsto e(x^{\sigma}\alpha y)$ is an alternator of $\mathcal{S}$. Hence $\mathcal{S}\subseteq\mathcal{A}_{b,\mathbb{D}}$. Yet $\dim\mathcal{A}_{b,\mathbb{D}}\leq d-1$ because $\mathcal{A}_{b,\mathbb{D}}$ is deeply intransitive, and we conclude that $\mathcal{S}=\mathcal{A}_{b,\mathbb{D}}$. Note here that $b:(x,y)\mapsto e(B(x,y))$ for the $\sigma$-sesquilinear form $B:(x,y)\mapsto x^{\sigma}\alpha y$. By point (c) of Proposition 5.4, we see that $b$ is nonisotropic.
From here, we deduce from point (b) of Proposition 5.4 that $\operatorname{\operatorname{Alt}}(\mathcal{S})$ has dimension $1$: it follows that for every nondegenerate bilinear form $b^{\prime}$ on $V$ such that $\mathcal{S}=\mathcal{A}_{b^{\prime},\mathbb{D}}$, we have $b^{\prime}\in\operatorname{\operatorname{Alt}}(\mathcal{S})$ and hence $b^{\prime}\in\mathbb{F}b$. This completes the proof of Theorem 1.10.
@@
where $b^{\prime\prime}:(x,y)\mapsto e(B^{\prime\prime}(x,y))$ for the $\sigma$-sesquilinear form $B^{\prime\prime}:(x,y)\mapsto B(\varphi(x),\varphi(y))$. Hence $\mathcal{A}_{b^{\prime\prime},\mathbb{D}}$ is the conjugate of $\mathcal{A}_{b,\mathbb{D}}$ under $\varphi^{-1}$, and more generally $\mathcal{A}_{\alpha b^{\prime\prime},\mathbb{D}}=\mathcal{A}_{b^{\prime\prime},\mathbb{D}}$ is conjugated to $\mathcal{A}_{b,\mathbb{D}}$ for all $\alpha\in\mathbb{F}^{\times}$.
Conversely, assume that $\mathcal{A}_{b,\mathbb{D}}$ is conjugated to $\mathcal{A}_{b^{\prime},\mathbb{D}}$ under an element of $\operatorname{\operatorname{GL}}_{\mathbb{D}}(V)$. Then by the above $\mathcal{A}_{b^{\prime},\mathbb{D}}=\mathcal{A}_{b^{\prime\prime},\mathbb{D}}$ for some $b^{\prime\prime}$ of the form $(x,y)\mapsto e(B^{\prime\prime}(x,y))$, where $B^{\prime\prime}$ is equivalent to $B$. The uniqueness statement in Theorem 1.10 yields $b^{\prime}=\alpha b^{\prime\prime}$ for some $\alpha\in\mathbb{F}^{\times}$, and hence $B^{\prime}=\alpha B^{\prime\prime}$. Hence $B^{\prime}$ is equivalent to $\alpha B$ for some $\alpha\in\mathbb{F}^{\times}$. This completes the proof of Theorem 1.11.
7 Semiaffine spaces of matrices with rank bounded below
In this short section, we apply the results on trivial spectrum subspaces to obtain significant ones on $\mathbb{F}$-affine subspaces of matrices with rank bounded below. This uses a strategy devised by Meshulam [10].
7.1 Semiaffine spaces of matrices with rank bounded below
@@
Let $n,p,r$ be integers with $\min(n,p)\geq r>0$. Let $\mathcal{S}$ be an $\mathbb{F}$-affine subspace of $\operatorname{\operatorname{M}}_{n,p}(\mathbb{D})$ in which every element has rank at least $r$. Assume that $|\mathbb{F}|\geq rd$. Then
$\operatorname{\operatorname{codim}}_{\operatorname{\operatorname{M}}_{n,p}(\mathbb{D})}\mathcal{S}\geq r+d\,\dbinom{r}{2}$,
where the codimension is to be understood as a codimension over $\mathbb{F}$. Moreover, this inequality is optimal, and equality holds only if $\mathcal{S}$ contains a rank $r$ matrix.
The proof requires the following weak version of Theorem 1.6:
Proposition 7.2.
Let $r\geq 1$ be such that $|\mathbb{F}|\geq rd$, and let $n\geq r$. Let $S$ be an $\mathbb{F}$-trivial spectrum linear subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$. Then $\operatorname{\operatorname{codim}}_{\operatorname{\operatorname{M}}_{n}(\mathbb{D})}S\geq r+d\dbinom{r}{2}$, with equality attained only if $n=r$.
For large values of $n$ with respect to $r$, this inequality can be considerably improved with limited effort, but we will not need such refinements.
Proof of Proposition 7.2.
We prove the result by induction on $n$. The case $n=r$ is a reformulation of Theorem 1.6. Assume now that $n>r$. Denote by $x$ the first vector of the standard basis of $\mathbb{D}^{n}$. Then $\dim_{\mathbb{F}}(Sx)\leq nd-1$ because $S$ is intransitive. Denote by $S^{\prime}$ the set of all matrices $N\in\operatorname{\operatorname{M}}_{n-1}(\mathbb{D})$ such that $\begin{bmatrix}0&[0]_{1\times(n-1)}\\ [0]_{(n-1)\times 1}&N\end{bmatrix}$ belongs to $S$, and note that $S^{\prime}$ is an $\mathbb{F}$-trivial spectrum linear subspace of $\operatorname{\operatorname{M}}_{n-1}(\mathbb{D})$. By the rank theorem
$\operatorname{\operatorname{codim}}_{\operatorname{\operatorname{M}}_{n}(\mathbb{D})}S\geq\operatorname{\operatorname{codim}}_{\mathbb{D}^{n}}(Sx)+\operatorname{\operatorname{codim}}_{\operatorname{\operatorname{M}}_{n-1}(\mathbb{D})}S^{\prime},$
and by induction $\operatorname{\operatorname{codim}}_{\operatorname{\operatorname{M}}_{n}(\mathbb{D})}S^{\prime}\geq r+d\dbinom{r}{2}$. Hence $\operatorname{\operatorname{codim}}_{\operatorname{\operatorname{M}}_{n}(\mathbb{D})}S\geq 1+r+d\dbinom{r}{2}$, which completes the proof. ∎
We can now prove Theorem 7.1.
Proof of Theorem 7.1.
The optimality of the inequality is exemplified by taking the space of all matrices of the form
$\begin{bmatrix}I_{r}+M&[?]_{r\times(p-r)}\\ [?]_{(n-r)\times r}&[?]_{(n-r)\times(p-r)}\end{bmatrix}$
where $M$ ranges over an $\mathbb{F}$-trivial spectrum subspace $\mathcal{M}$ of $\operatorname{\operatorname{M}}_{r}(\mathbb{F})$ with dimension $\alpha(r,d)$ (say, one that is described on page 1.1). It is clear that the codimension of this space in $\operatorname{\operatorname{M}}_{n,p}(\mathbb{D})$ is the one of $\mathcal{M}$ in $\operatorname{\operatorname{M}}_{r}(\mathbb{D})$, that is $r+d\,\dbinom{r}{2}$. Every matrix in that space has its upper-left $r\times r$ block invertible, and hence its rank is at least $r$.
Now, denote by $s$ the least possible rank in $\mathcal{S}$ (at this point it may very well be that $s>r$). Replacing $\mathcal{S}$ with an equivalent space if necessary, we can then assume that $\mathcal{S}$ contains $J_{s}=\begin{bmatrix}I_{s}&[0]_{s\times(p-s)}\\ [0]_{(n-s)\times s}&[0]_{(n-s)\times(p-s)}\end{bmatrix}$. Denote by $\overrightarrow{\mathcal{S}}$ the translation vector space of $\mathcal{S}$, and write every matrix $M$ of it as $M=\begin{bmatrix}A(M)&C(M)\\ B(M)&D(M)\end{bmatrix}$ with $A(M)\in\operatorname{\operatorname{M}}_{s}(\mathbb{F})$ and so on. Consider the $\mathbb{F}$-linear subspace $W$ of $\overrightarrow{\mathcal{S}}$ consisting of all the matrices $M\in\overrightarrow{\mathcal{S}}$ such that $B(M)=0$, $C(M)=0$ and $D(M)=0$. Hence $\operatorname{\operatorname{codim}}_{\operatorname{\operatorname{M}}_{n,p}(\mathbb{D})}\mathcal{S}\geq\operatorname{\operatorname{codim}}_{\operatorname{\operatorname{M}}_{s}(\mathbb{D})}A(W)$. For all $M\in W$, the matrix
$J_{s}+M=\begin{bmatrix}I_{s}+A(M)&[0]_{s\times(p-s)}\\ [0]_{(n-s)\times s}&[0]_{(n-s)\times(p-s)}\end{bmatrix}$
has rank at least $s$ because it belongs to $\mathcal{S}$, and hence $I_{s}+A(M)$ is invertible. It follows that $A(W)$ is an $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{M}}_{s}(\mathbb{D})$. Since $|\mathbb{F}|\geq rd\geq sd$, we deduce from Proposition 7.2 that $\operatorname{\operatorname{codim}}_{\operatorname{\operatorname{M}}_{s}(\mathbb{D})}A(W)\geq r+d\dbinom{r}{2}$, with strict equality if $s>r$, and we conclude that $\operatorname{\operatorname{codim}}_{\operatorname{\operatorname{M}}_{n,p}(\mathbb{D})}\mathcal{S}\geq r+d\dbinom{r}{2}$, with strict inequality if $s>r$. This proves all the claimed statements. ∎
Remark 7.1.
If we had used Theorem 1.6 instead of Proposition 7.2, we would have needed the inequality $|\mathbb{F}|\geq d\max(n,p)$ because we have no prior control on the least possible rank in $\mathcal{S}$. Clearly, this requirement is not very satisfying, hence our insistence in requiring only that $|\mathbb{F}|\geq rd$.
7.2 Optimal semiaffine spaces of nonsingular matrices
At this point, it is natural that we wish to obtain a description of the optimal $\mathbb{F}$-affine subspaces of matrices with rank at least $r$, meaning the ones with the minimal possible codimension in $\operatorname{\operatorname{M}}_{n,p}(\mathbb{D})$. Judging from the length of [16], which covers only the case $\mathbb{D}=\mathbb{F}$, it is safer to leave this problem for future work. Nevertheless, we will tackle here the limit case $n=p=r$ as it is can be obtained with few technical difficulties from our results on trivial spectrum linear subspaces.
Let us start with some terminology and notation.
Definition 7.1.
Two subsets $\mathcal{X}$ and $\mathcal{Y}$ of $\operatorname{\operatorname{M}}_{n,p}(\mathbb{D})$ are called equivalent , and we write $\mathcal{X}\sim\mathcal{Y}$, when there exists a pair $(P,Q)\in\operatorname{\operatorname{GL}}_{n}(\mathbb{D})\times\operatorname{\operatorname{GL}}_{p}(\mathbb{D})$ such that $\mathcal{Y}=P\mathcal{X}Q$, i.e., when $\mathcal{X}$ and $\mathcal{Y}$ represent the same set of linear mappings from a $p$-dimensional $\mathbb{D}$-vector space to an $n$-dimensional $\mathbb{D}$-vector space, in a different choice of bases of these spaces.
@@ 7.2 Optimal semiaffine spaces of nonsingular matrices
Let $n\geq 1$. Assume that $\mathbb{D}$ is not of quadratic type over $\mathbb{F}$, and that $|\mathbb{F}|\geq nd$. Then, up to equivalence, there is exactly one optimal $\mathbb{F}$-affine subspace of invertible matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$. One such space is the $n$-fold joint $(1_{\mathbb{D}}+H)\vee\cdots\vee(1_{\mathbb{D}}+H)$, where $H$ is an arbitrary $\mathbb{F}$-linear hyperplane of $\mathbb{D}$ that does not contain $1_{\mathbb{D}}$.
@@
Let $\mathcal{M}$ be an optimal $\mathbb{F}$-affine subspace of invertible matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$. Then there exists a partition $n=n_{1}+\cdots+n_{p}$ and a list $(P_{1},\dots,P_{p})$ of $e$-nonisotropic matrices, with $P_{i}\in\operatorname{\operatorname{GL}}_{n_{i}}(\mathbb{D})$, such that
@@
Conversely, let $n=n_{1}+\cdots+n_{p}$ be a partition, and $(P_{1},\dots,P_{p})$ be a list of $e$-nonisotropic matrices, with $P_{i}\in\operatorname{\operatorname{GL}}_{n_{i}}(\mathbb{D})$. Then $(P_{1}+\mathcal{S}\mathcal{H}_{n_{1}}(\mathbb{D}))\vee\cdots\vee(P_{p}+\mathcal{S}\mathcal{H}_{n_{p}}(\mathbb{D}))$ is an optimal $\mathbb{F}$-affine subspace of invertible matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$.
@@
Let $\mathcal{S}$ be an irreducible optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, where $V$ is an $n$-dimensional vector space over $\mathbb{D}$, with $|\mathbb{F}|\geq nd$. Then $\mathbb{F}x+\mathcal{S}x=V$ for all $x\in V\smallsetminus\{0\}$.
@@
We combine Theorems 1.8 and 1.10 to find that $\mathcal{S}=\mathcal{A}_{b,\mathbb{D}}$ for some nondegenerate $\mathbb{F}$-bilinear form $b$ of type $(x,y)\mapsto e(B(x,y))$, where $B$ is a nondegenerate $\sigma$-sesquilinear form on $V$. Let $x\in V\smallsetminus\{0\}$. By point (b) of Proposition 5.4 we find $\dim_{\mathbb{F}}(\mathcal{S}x)=nd-1$. The conclusion is obtained by noting that $x\not\in\mathcal{S}x$ because $\mathcal{S}$ has $\mathbb{F}$-trivial spectrum. ∎
@@
The second statement is obvious because, for every $k\in\mathopen{[\![}1,p\mathclose{]\!]}$ and every $e$-nonisotropic $P\in\operatorname{\operatorname{GL}}_{k}(\mathbb{D})$ we have $P+\mathcal{M}_{k}\sim I_{k}+P^{-1}\mathcal{M}_{k}$, and $P^{-1}\mathcal{M}_{k}$ has $\mathbb{F}$-trivial spectrum and dimension $\alpha(k,d)$ over $\mathbb{F}$; note also, for the remainder of the proof, that $\operatorname{\operatorname{span}}_{\mathbb{F}}(P+\mathcal{M}_{k})X=\mathbb{D}^{k}$ for all $X\in\mathbb{D}^{k}\smallsetminus\{0\}$, as a consequence of Lemma 7.5.
It remains to prove the uniqueness statement. It is better to think in geometrical terms here. So, let $V$ be an $n$-dimensional vector space over $\mathbb{D}$, and let $\mathcal{S}$ be an optimal $\mathbb{F}$-affine subspace of automorphisms of $V$. Assume that we have bases $(e_{1},\dots,e_{n})$ and $(f_{1},\dots,f_{n})$ of $V$ such that the matrices that represent the elements of $\mathcal{S}$ form the space $(P_{1}+\mathcal{M}_{n_{1}})\vee\cdots\vee(P_{p}+\mathcal{M}_{n_{p}})$. We consider the integers $N_{i}:=\sum_{k=1}^{i}n_{k}$ and the spaces $U_{i}:=\operatorname{\operatorname{span}}(e_{1},\dots,e_{N_{i}})$ and $V_{i}:=\operatorname{\operatorname{span}}(f_{1},\dots,f_{N_{i}})$ for $i\in\mathopen{[\![}0,p\mathclose{]\!]}$. We shall prove that any other choice of bases will yield the same spaces. To see this, we look at the dimension $\dim_{\mathbb{F}}(\mathcal{S}x)$ when $x$ ranges over $V$. Using the previous remark, we see that for all $i\in\mathopen{[\![}1,p\mathclose{]\!]}$ and all $x\in U_{i}\smallsetminus U_{i-1}$, we have $\operatorname{\operatorname{span}}_{\mathbb{F}}(\mathcal{S}x)=V_{i}$. Hence the spaces $V_{0},\dots,V_{p}$ are uniquely determined by $\mathcal{S}$ only, as the possible $\mathcal{S}x$ spaces when $x$ ranges over $V$; and then $U_{i}=\{x\in V:\;\mathcal{S}x\subseteq V_{i}\}$ is also uniquely determined by $\mathcal{S}$ for all $i\in\mathopen{[\![}1,p\mathclose{]\!]}$.

### v2

@@ 6.3 The full structure theorem for optimal spaces
Let $V$ be a (right) $\mathbb{D}$-vector space with finite dimension $n>0$. Assume that $|\mathbb{F}|\geq nd$, and let $b$ be a nonisotropic $\mathbb{F}$-bilinear form on $V$ such that $S:=\mathcal{A}_{b,\mathbb{D}}$ is an optimal irreducible trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$.
We shall prove that $\dim\operatorname{\operatorname{Alt}}(S)=1$ in any case.
  • Assume first that $n\geq\max(4-d,2)$. Then $\dim S=\alpha(n,d)\geq\alpha(n,d)-n+\max(4-d,2)$ and the claimed statement is deduced from point (c) of Theorem 1.22.
  • Assume now that $n=2$ and $d=1$. Then $\dim S=1$, and we take $u_{0}\in S\smallsetminus\{0\}$. Since $S$ is irreducible, $u_{0}$ must be invertible, and we deduce that $\dim Sx=1=nd-1$ for all $x\in V\smallsetminus\{0\}$. Moreover $S$ is target-reduced because $u_{0}$ is surjective. Hence Lemma 5.3 shows that $\dim\operatorname{\operatorname{Alt}}(S)\leq 1$, and we conclude that $\dim\operatorname{\operatorname{Alt}}(S)=1$ because $b\in\operatorname{\operatorname{Alt}}(S)$.
  • If $n=1$ and $d>1$ then $\dim_{\mathbb{F}}(Sx)=\dim S=nd-1$ for all $x\in V\smallsetminus\{0\}$, and obviously $S$ is target-reduced, so once more we find by Proposition 5.3 that $\dim\operatorname{\operatorname{Alt}}(S)=1$.
  • If $n=1$ and $d=1$ then $S=\{0\}$ and $\operatorname{\operatorname{Alt}}(S)$ is the space of all $\mathbb{F}$-bilinear forms on $V$, which has dimension $1$.
Then, we apply Proposition 5.2 and find that $\mathbb{D}$ has quadratic type over $\mathbb{F}$ and that, for an associated pair $(\sigma,e)$, there exists a $\sigma$-sesquilinear form $B:V^{2}\rightarrow\mathbb{D}$ such that $b:(x,y)\mapsto e(B(x,y))$. Besides, it has already been observed in Section 2 that $B$ is uniquely determined by $b$.
@@ 6.4 The quadratic type case
We finish this section by proving the results that are specific to the division algebras of quadratic type: Theorems 1.12 and 1.13.
So, assume that $\mathbb{D}$ has quadratic type over $\mathbb{F}$, and denote by $(\sigma,e)$ an associated pair. Let $V$ be an $n$-dimensional vector space over $\mathbb{D}$, with $n\geq 1$. Let $S$ be an irreducible optimal trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ such that $|\mathbb{F}|\geq nd$.
If $n\geq 2$ then we apply Theorems 1.9 and 1.10 to find a $\sigma$-sesquilinear nondegenerate form $B:V^{2}\rightarrow\mathbb{D}$ such that $S=\mathcal{A}_{b,\mathbb{D}}$ for $b:(x,y)\mapsto e(B(x,y))$.
Assume now that $n=1$. In that case we need only consider the special case where $V=\mathbb{D}$. By the isomorphism $a\in\mathbb{D}\mapsto[x\mapsto ax]\in\operatorname{\operatorname{End}}_{\mathbb{D}}(\mathbb{D})$ of $\mathbb{F}$-vector spaces, we associate to $S$ an $\mathbb{F}$-linear hyperplane $H$ of $\mathbb{D}$. Thus $H$ is the kernel of $x\mapsto e(\alpha x)$ for some $\alpha\in\mathbb{D}^{\times}$. Then, for all $a\in H$, we have
@@
This shows that the nondegenerate bilinear form $b:(x,y)\in\mathbb{D}^{2}\mapsto e(x^{\sigma}\alpha y)$ is an alternator of $S$. Hence $S\subseteq\mathcal{A}_{b,\mathbb{D}}$. Yet $\dim\mathcal{A}_{b,\mathbb{D}}\leq d-1$ because $\mathcal{A}_{b,\mathbb{D}}$ is deeply intransitive, and we conclude that $S=\mathcal{A}_{b,\mathbb{D}}$. Note here that $b:(x,y)\mapsto e(B(x,y))$ for the $\sigma$-sesquilinear form $B:(x,y)\mapsto x^{\sigma}\alpha y$. By point (c) of Proposition 5.4, we see that $b$ is nonisotropic.
From here, we deduce from point (b) of Proposition 5.4 that $\operatorname{\operatorname{Alt}}(S)$ has dimension $1$: it follows that for every nondegenerate bilinear form $b^{\prime}$ on $V$ such that $S=\mathcal{A}_{b^{\prime},\mathbb{D}}$, we have $b^{\prime}\in\operatorname{\operatorname{Alt}}(S)$ and hence $b^{\prime}\in\mathbb{F}b$. This completes the proof of Theorem 1.12.
@@
where $b^{\prime\prime}:(x,y)\mapsto e(B^{\prime\prime}(x,y))$ for the $\sigma$-sesquilinear form $B^{\prime\prime}:(x,y)\mapsto B(\varphi(x),\varphi(y))$. Hence $\mathcal{A}_{b^{\prime\prime},\mathbb{D}}$ is the conjugate of $\mathcal{A}_{b,\mathbb{D}}$ under $\varphi^{-1}$, and more generally $\mathcal{A}_{\alpha b^{\prime\prime},\mathbb{D}}=\mathcal{A}_{b^{\prime\prime},\mathbb{D}}$ is similar to $\mathcal{A}_{b,\mathbb{D}}$ for all $\alpha\in\mathbb{F}^{\times}$.
Conversely, assume that $\mathcal{A}_{b,\mathbb{D}}$ is similar to $\mathcal{A}_{b^{\prime},\mathbb{D}}$ under an element of $\operatorname{\operatorname{GL}}_{\mathbb{D}}(V)$. Then by the above $\mathcal{A}_{b^{\prime},\mathbb{D}}=\mathcal{A}_{b^{\prime\prime},\mathbb{D}}$ for some $b^{\prime\prime}$ of the form $(x,y)\mapsto e(B^{\prime\prime}(x,y))$, where $B^{\prime\prime}$ is equivalent to $B$. The uniqueness statement in Theorem 1.12 yields $b^{\prime}=\alpha b^{\prime\prime}$ for some $\alpha\in\mathbb{F}^{\times}$, and hence $B^{\prime}=\alpha B^{\prime\prime}$. Hence $B^{\prime}$ is equivalent to $\alpha B$ for some $\alpha\in\mathbb{F}^{\times}$. This completes the proof of Theorem 1.13.
7 Optimal affine spaces of units
In this final section, we solve our initial problem of unital affine subspaces of a simple algebra.
7.1 The greatest possible dimension
To start with, we have a straightforward application of Theorem 1.8:
@@ 7.1 The greatest possible dimension
Let $\mathbb{D}$ be a division ring with finite dimension $d$ over a central subfield $\mathbb{F}$, and let $n\geq 1$ be such that $|\mathbb{F}|\geq nd$. Then the greatest possible dimension for a unital $\mathbb{F}$-affine subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ is $n(d-1)+d\binom{n}{2}$.
7.2 The optimal spaces
For the structure of optimal spaces, we will split the discussion, whether $\mathbb{D}$ has quadratic type over $\mathbb{F}$ or not.
@@ 7.2 The optimal spaces
Let $n\geq 1$. Assume that $\mathbb{D}$ is not of quadratic type over $\mathbb{F}$, and that $|\mathbb{F}|\geq nd$. Then, up to equivalence, there is exactly one optimal unital $\mathbb{F}$-affine subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$.
One such space is the $n$-fold joint $(1_{\mathbb{D}}+H)\vee\cdots\vee(1_{\mathbb{D}}+H)$, where $H$ is an arbitrary $\mathbb{F}$-linear hyperplane of $\mathbb{D}$ that does not contain $1_{\mathbb{D}}$.
@@
Let $\mathcal{M}$ be an optimal unital $\mathbb{F}$-affine subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$. Then there exists a partition $n=n_{1}+\cdots+n_{p}$ and a list $(P_{1},\dots,P_{p})$ of $e$-nonisotropic matrices, with $P_{i}\in\operatorname{\operatorname{GL}}_{n_{i}}(\mathbb{D})$, such that
@@
Conversely, let $n=n_{1}+\cdots+n_{p}$ be a partition, and $(P_{1},\dots,P_{p})$ be a list of $e$-nonisotropic matrices, with $P_{i}\in\operatorname{\operatorname{GL}}_{n_{i}}(\mathbb{D})$. Then $(P_{1}+\mathcal{S}\mathcal{H}_{n_{1}}(\mathbb{D}))\vee\cdots\vee(P_{p}+\mathcal{S}\mathcal{H}_{n_{p}}(\mathbb{D}))$ is an optimal unital $\mathbb{F}$-affine subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$.
@@
Let $S$ be an irreducible optimal trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, where $V$ is an $n$-dimensional vector space over $\mathbb{D}$, with $|\mathbb{F}|\geq nd$. Then $\mathbb{F}x+Sx=V$ for all $x\in V\smallsetminus\{0\}$.
@@
We combine Theorems 1.10 and 1.12 to find that $S=\mathcal{A}_{b,\mathbb{D}}$ for some nondegenerate $\mathbb{F}$-bilinear form $b$ of type $(x,y)\mapsto e(B(x,y))$, where $B$ is a nondegenerate $\sigma$-sesquilinear form on $V$. Let $x\in V\smallsetminus\{0\}$. By point (b) of Proposition 5.4 we find $\dim_{\mathbb{F}}(Sx)=nd-1$. The conclusion is obtained by noting that $x\not\in Sx$ because $S$ has trivial spectrum. ∎
@@
The second statement is obvious because, for every $k\in\mathopen{[\![}1,p\mathclose{]\!]}$ and every $e$-nonisotropic $P\in\operatorname{\operatorname{GL}}_{k}(\mathbb{D})$ we have $P+\mathcal{M}_{k}\sim I_{k}+P^{-1}\mathcal{M}_{k}$, and $P^{-1}\mathcal{M}_{k}$ has trivial spectrum and dimension $\alpha(k,d)$ over $\mathbb{F}$; note also, for the remainder of the proof, that $\operatorname{\operatorname{span}}_{\mathbb{F}}(P+\mathcal{M}_{k})X=\mathbb{D}^{k}$ for all $X\in\mathbb{D}^{k}\smallsetminus\{0\}$, as a consequence of Lemma 7.4.
It remains to prove the uniqueness statement. It is better to think in geometrical terms here. So, let $V$ be an $n$-dimensional vector space over $\mathbb{D}$, and let $S$ be an optimal $\mathbb{F}$-affine subspace of automorphisms of $V$. Assume that we have bases $(e_{1},\dots,e_{n})$ and $(f_{1},\dots,f_{n})$ of $V$ such that the matrices that represent the elements of $S$ form the space $(P_{1}+\mathcal{M}_{n_{1}})\vee\cdots\vee(P_{p}+\mathcal{M}_{n_{p}})$. We consider the integers $N_{i}:=\sum_{k=1}^{i}n_{k}$ and the spaces $U_{i}:=\operatorname{\operatorname{span}}(e_{1},\dots,e_{N_{i}})$ and $V_{i}:=\operatorname{\operatorname{span}}(f_{1},\dots,f_{N_{i}})$ for $i\in\mathopen{[\![}0,p\mathclose{]\!]}$. We shall prove that any other choice of bases will yield the same spaces. To see this, we look at the dimension $\dim_{\mathbb{F}}(Sx)$ when $x$ ranges over $V$. Using the previous remark, we see that for all $i\in\mathopen{[\![}1,p\mathclose{]\!]}$ and all $x\in U_{i}\smallsetminus U_{i-1}$, we have $\operatorname{\operatorname{span}}_{\mathbb{F}}(Sx)=V_{i}$. Hence the spaces $V_{0},\dots,V_{p}$ are uniquely determined by $S$ only, as the possible $Sx$ spaces when $x$ ranges over $V$; and then $U_{i}=\{x\in V:\;Sx\subseteq V_{i}\}$ is also uniquely determined by $S$ for all $i\in\mathopen{[\![}1,p\mathclose{]\!]}$.